let A(2,-3) and B(-2,1) be vertices of triangle ABC,if the centroid of the triangle moves on the line 2x+3y=1 then the locus of the vertex C is
 

Grade:10

1 Answers

Rishita Agrawal
5 years ago
Let C(h,k)
To find the centroid, (x+x+x3/3, y1  + y+y3/3)
(2-2+h/3, -3+1+k/3)
(h/3, k-2/3)
When we replace these values in equation of centroid,
​2.h/3 + 3. (k-2)/3 = 1
2h+3k=9
Putting x as h and y as k,
2x+3y=9